OpenStudy (anonymous):

Alright they give me a right triangle and they give me a interior angle of 58 and a outside of 4.0 I need help on how to solve this

OpenStudy (unklerhaukus):

can you draw it, with the draw button

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

what is 4.0? a side-length?

OpenStudy (anonymous):



OpenStudy (anonymous):

thats how it looks

OpenStudy (anonymous):

and they want me to solve for x

OpenStudy (anonymous):

yes it is a side length

OpenStudy (anonymous):

how would i solve this ?

OpenStudy (unklerhaukus):

do you remember that \[\tan\theta=\frac{\text{opposite side}}{\text{adjacent side}}\]?

OpenStudy (anonymous):

i dont really remember it how would i plug that in with my question ?

OpenStudy (unklerhaukus):

well you have theta is 58°, and you have the opposite side is 4.0

OpenStudy (unklerhaukus):

\[{\text{adjacent side}}=\frac{\text{opposite side}}{\tan\theta}\]

OpenStudy (unklerhaukus):

x = 4.0 / tan(58°)

OpenStudy (anonymous):

i got 14.5 i dont know if thats right